A 0.50-mm-wide slit is illuminated by light of wavelength 500 nm. What is the width of the central maximum on a screen 2.0 m behind the slit? Please answer in mm.
Answer #1

given

d = 0.5 mm = 0.5*10^-3 m

lamda = 500 nm = 500*10^-9 m

R = 2.0 m

we know, widht of central maximum, w = 2*lamda*R/d

= 2*500*10^-9*2/(0.5*10^-3)

= 0.004 m

= 4.0 mm